OpenStudy (anonymous):

I WILL GIVE MEDAL IF U HELP Use properties to evaluate the expression. (13 + 6) + 4 A. 19 B. 23 C. 27 D. 37

OpenStudy (anonymous):

B is the answer, though--- PEMDAS says so.

OpenStudy (anonymous):

:D thx ok ill remember pemdas i forgot about that XD

OpenStudy (anonymous):

Thank you and if you need a reminder: Parenthesis Exponents Multiplication Division Addiotion and Subtraction.

OpenStudy (anonymous):

:D thx

OpenStudy (wolfmegan2612):

13+6=19+4=23

OpenStudy (anonymous):

i got the help i needed ima close this now :D

OpenStudy (anonymous):

thanks to all ppl